What is the energy of the photon that, when absorb
Answer #1

1)energy of photon E=-13.6z2(1/n12-1/n22)

here z=1,n1=1,n2=4

E=-13.6(1/12-1/42)=-12.75 ev

2)here z=1,n1=4,n2=8

energy E=-13.6(1/42-1/82)=-0.6375 ev